40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 174     Entry time: Wed Oct 23 02:45:07 2013
Author: Koji 
Type: General 
Category: General 
Subject: PD realignment 

DCPD2 got misaligned during the cable installation. The PD alignment procedure have been gone through again.

Cavity locking

- Removed the FC layers for the cavity related mirrors.

- Aligned and locked the cavity.

PD alignment

- Loosen DCPD2. Checked the reflection with a IR card. Checked the spot on the PD with an IR viewer.

- Finger-tight the screws. Check the reflection with the card again. Check the pot on the PD with a CCD.

- If the spot positions are not satisfactory repeat the process.

- If the spot positions are satisfactory, take pictures of the CCD image.

- Fixing screws for all of the PDs/QPDs were tighten by the torque driver with a torque od 1.75 inch lb.

PD QE measurements

- Measure the power incident on the PDs.

- Set up the transimpedance amp to check the photo current.

- PD1 (T side) 9.10+/-0.03 V 13.02 +/- 0.01W -> QE ~80%

- PD2 (R side) 8.70+/-0.01 V 12.53 +/- 0.01W -> QE ~80%

- These are not strange values considering the presence of the glass caps.

PZT polarity check

- The connections between the PZT electrodes and the pins were checked.

- The positive side is marked by a knot on the wire.

FC painting

- The new FC bottle was brought from Downs, thanks to Margot.

 

ELOG V3.1.3-